OpenStudy (anonymous):

How do you find a domain? for example: let f(x)=x+5 and g(x)=x^2-25. What is the domain of f/g (x)?

OpenStudy (solomonzelman):

The function will be x+5 ---- x²-25 factoring the bottom, and canceling x+5, f/g (x) = 1/(x-5)

OpenStudy (solomonzelman):

The domain can be any thing, but the restriction is 5.

OpenStudy (solomonzelman):

(-∞,5) Union (5,∞)

OpenStudy (anonymous):

how would you factor the bottom

OpenStudy (solomonzelman):

x²-25 = x² - 5² = (x-5)(x+5)

OpenStudy (solomonzelman):

clear with that ?

OpenStudy (solomonzelman):

I have finally started on limits officially, and this really reminds me one of those algebraically easy ones.

OpenStudy (aum):

The domain should be found BEFORE simplifying f/g(x). Therefore, domain is all real numbers except x = -5 and x = +5.

OpenStudy (anonymous):

yes thank you

OpenStudy (aum):

In case you need confirmation: http://www.wolframalpha.com/input/?i=domain+of+%28x%2B5%29%2F%28x^2-25%29

OpenStudy (solomonzelman):

aum, I thought you cancel the x+5, don't you ?

OpenStudy (aum):

You can simplify the expression. But when domain is asked for a rational expression you need to find what makes the denominator zero BEFORE simplifying.

OpenStudy (solomonzelman):

well, all I can tell you.... \(\LARGE\color{black}{ \lim_{x \rightarrow C} ~\frac{x+5}{x^2-25} }\) \(\LARGE\color{black}{ \lim_{x \rightarrow C} ~\frac{x+5}{(x+5)(x-5)} }\) \(\LARGE\color{black}{ \lim_{x \rightarrow C} ~x+5 }\) so limit x→5 and f(5), (I think) both exist)

OpenStudy (solomonzelman):

wait, no they don't nvm

OpenStudy (solomonzelman):

The limit exists, and the limit as x→C of f(x) = f(C)

OpenStudy (solomonzelman):

and the sides are equal..

OpenStudy (solomonzelman):

BUT f(C) does not exist.

OpenStudy (aum):

The function is continuous at x = -5. No dispute there. But when f(x) is divided by g(x) then all values that make g(x) zero have to be excluded from the domain.

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \lim_{x \rightarrow -5} ~\frac{x+5}{x^2-25} }\) I like more, \(\LARGE\color{black}{ \lim_{x \rightarrow -5} ~\frac{\frac{d}{dx}(x+5)}{\frac{d}{dx}(x^2-25)} }\) \(\LARGE\color{black}{ \lim_{x \rightarrow -5} ~\frac{1}{2x} }\) \(\LARGE\color{black}{ \frac{1}{-10} }\)

OpenStudy (solomonzelman):

continuous at x=... doesn't tend to make sense to me though, but I can somewhat see what you are saying.

OpenStudy (solomonzelman):

L'H'S ?! just showing off.

OpenStudy (solomonzelman):

Continuous function has to have 3 conditions. f(c) exists lim x→c f(c) exist lim x→c f(x) = f(c)